\begin{thebibliography}{99}


%\cite{Vasiliev:2004qz}
\bibitem{Vasiliev:2004qz}
   M.~A.~Vasiliev,
   %``Higher spin gauge theories in various dimensions,''
   Fortsch.\ Phys.\  {\bf 52}, 702 (2004)
   [arXiv:hep-th/0401177].
%%CITATION = HEP-TH 0401177;%%
%\cite{Bekaert:2005vh}
%\bibitem{Bekaert:2005vh}
   X.~Bekaert, S.~Cnockaert, C.~Iazeolla and M.~A.~Vasiliev,
   %``Nonlinear higher spin theories in various dimensions,''
   [arXiv:hep-th/0503128].
   %%CITATION = HEP-TH 0503128;%%
%\cite{Sorokin:2004ie}
%\bibitem{Sorokin:2004ie}
   D.~Sorokin,
   %``Introduction to the classical theory of higher spins,''
   AIP Conf.\ Proc.\  {\bf 767}, 172 (2005)
   [arXiv:hep-th/0405069].
   %%CITATION = HEP-TH 0405069;%%
   %\cite{Bouatta:2004kk}
%\bibitem{Bouatta:2004kk}
   N.~Bouatta, G.~Compere and A.~Sagnotti,
   %``An introduction to free higher-spin fields,''
   [arXiv:hep-th/0409068].
   %%CITATION = HEP-TH 0409068;%%
   %\cite{Francia:2006hp}
%\cite{Bekaert:2003uc}
%\bibitem{Bekaert:2003uc}
   X.~Bekaert, I.~L.~Buchbinder, A.~Pashnev and M.~Tsulaia,
   %``On higher spin theory: Strings, BRST, dimensional reductions,''
   Class.\ Quant.\ Grav.\  {\bf 21} (2004) S1457
   [arXiv:hep-th/0312252].
   %%CITATION = HEP-TH 0312252;%%
%\cite{Campoleoni:2009je}
%\bibitem{Campoleoni:2009je}
  A.~Campoleoni,
  %``Metric-like Lagrangian Formulations for Higher-Spin Fields of Mixed
  %Symmetry,''
  arXiv:0910.3155,
  %%CITATION = ARXIV:0910.3155;%%
%\cite{Francia:2010ap}
%\bibitem{Francia:2010ap}
  D.~Francia,
  %``On the relation between local and geometric Lagrangians for higher spins,''
  J.\ Phys.\ Conf.\ Ser.\  {\bf 222}, 012002 (2010)
  [arXiv:1001.3854].
  %%CITATION = 00462,222,012002;%%
%\cite{Bekaert:2010hw}
%\bibitem{Bekaert:2010hw}
  X.~Bekaert, N.~Boulanger and P.~Sundell,
  %``How higher-spin gravity surpasses the spin two barrier: no-go theorems
  %versus yes-go examples,''
  [arXiv:1007.0435].
  %%CITATION = ARXIV:1007.0435;%%



%\cite{Fotopoulos:2008ka}
\bibitem{Fotopoulos:2008ka}
  A.~Fotopoulos and M.~Tsulaia,
  %``Gauge Invariant Lagrangians for Free and Interacting Higher Spin Fields. A
  %Review of the BRST formulation,''
  Int.\ J.\ Mod.\ Phys.\  A {\bf 24}, 1 (2009)
  [arXiv:0805.1346 [hep-th]].
  %%CITATION = IMPAE,A24,1;%%


%\cite{Fradkin:1986qy}
\bibitem{Fradkin:1986qy}
   E.~S.~Fradkin and M.~A.~Vasiliev,
   %``Cubic Interaction In Extended Theories Of Massless Higher Spin Fields,''
   Nucl.\ Phys.\ B {\bf 291}, 141 (1987),
   %%CITATION = NUPHA,B291,141;%%
%\cite{Fradkin:1986ka}
%\bibitem{Fradkin:1986ka}
%   E.~S.~Fradkin and M.~A.~Vasiliev,
   %``Candidate To The Role Of Higher Spin Symmetry,''
   Annals Phys.\  {\bf 177}, 63 (1987),
   %%CITATION = APNYA,177,63;%%
%\cite{Vasiliev:1990en}
%\bibitem{Vasiliev:1990en}
   M.~A.~Vasiliev,
   %``Consistent Equation For Interacting Gauge Fields Of All Spins In
   %(3+1)-Dimensions,''
   Phys.\ Lett.\ B {\bf 243}, 378 (1990),
   %%CITATION = PHLTA,B243,378;%%
   %\cite{Vasiliev:2001wa}
  %%CITATION = HEP-TH 0106200;%%
   %\cite{Vasiliev:1992av}
%\bibitem{Vasiliev:1992av}
  % M.~A.~Vasiliev,
   %``More on equations of motion for interacting massless fields of all spins in
   %(3+1)-dimensions,''
   Phys.\ Lett.\ B {\bf 285}, 225 (1992),
   %%CITATION = PHLTA,B285,225;%%
%\cite{Vasiliev:2003ev}
%\bibitem{Vasiliev:2003ev}
  % M.~A.~Vasiliev,
   %``Nonlinear equations for symmetric massless higher spin fields in
   %(A)dS(d),''
   Phys.\ Lett.\ B {\bf 567}, 139 (2003)
   [arXiv:hep-th/0304049],
   %%CITATION = HEP-TH 0304049;%%
%\cite{Vasiliev:2007yc}
%\bibitem{Vasiliev:2007yc}
 % M.~A.~Vasiliev,
  %``On Conformal, SL(4,R) and Sp(8,R) Symmetries of 4d Massless Fields,''
  arXiv:0707.1085 [hep-th],
  %%CITATION = ARXIV:0707.1085;%%
%\cite{Vasiliev:2004cm}
%\bibitem{Vasiliev:2004cm}
 % M.~A.~Vasiliev,
  %``Higher spin superalgebras in any dimension and their representations,''
  JHEP {\bf 0412} (2004) 046
  [arXiv:hep-th/0404124],
  %%CITATION = JHEPA,0412,046;%%
%\cite{Skvortsov:2006at}
%\bibitem{Skvortsov:2006at}
  E.~D.~Skvortsov and M.~A.~Vasiliev,
  %``Geometric formulation for partially massless fields,''
  Nucl.\ Phys.\  B {\bf 756} (2006) 117
  [arXiv:hep-th/0601095],
  %%CITATION = NUPHA,B756,117;%%
%\cite{Sorokin:2008tf}
%\bibitem{Sorokin:2008tf}
  D.~P.~Sorokin and M.~A.~Vasiliev,
  %``Reducible higher-spin multiplets in flat and AdS spaces and their geometric
  %frame-like formulation,''
  Nucl.\ Phys.\  B {\bf 809}, 110 (2009)
  [arXiv:0807.0206 [hep-th]],
  %%CITATION = NUPHA,B809,110;%%


%\cite{Bandos:1999qf}
\bibitem{Bandos:1999qf}
  I.~A.~Bandos, J.~Lukierski and D.~P.~Sorokin,
  %``Superparticle models with tensorial central charges,''
  Phys.\ Rev.\  D {\bf 61}, 045002 (2000)
  [arXiv:hep-th/9904109].
  %%CITATION = PHRVA,D61,045002;%%
%\cite{Vasiliev:2001zy}
%\bibitem{Vasiliev:2001zy}
  M.~A.~Vasiliev,
  %``Conformal higher spin symmetries of 4D massless supermultiplets and
  %osp(L,2M) invariant equations in generalized (super)space,''
  Phys.\ Rev.\  D {\bf 66}, 066006 (2002)
  [arXiv:hep-th/0106149].
  %%CITATION = PHRVA,D66,066006;%%





%\cite{Fronsdal:1978rb}
\bibitem{Fronsdal:1978rb}
  C.~Fronsdal,
  %``Massless Fields With Integer Spin,''
  Phys.\ Rev.\  D {\bf 18}, 3624 (1978),
  %%CITATION = PHRVA,D18,3624;%%
%\cite{Berends:1985xx}
%\bibitem{Berends:1985xx}
  F.~A.~Berends, G.~J.~H.~Burgers and H.~van Dam,
  %``Explicit Construction Of Conserved Currents For Massless Fields Of
  %Arbitrary Spin,''
  Nucl.\ Phys.\  B {\bf 271}, 429 (1986),
  %%CITATION = NUPHA,B271,429;%%
%\cite{Ouvry:1986dv}
%\bibitem{Ouvry:1986dv}
  S.~Ouvry and J.~Stern,
  %``Gauge Fields Of Any Spin And Symmetry,''
  Phys.\ Lett.\  B {\bf 177}, 335 (1986),
  %%CITATION = PHLTA,B177,335;%%
%\cite{Bengtsson:1984si}
%\cite{Bengtsson:1983pg}
%\bibitem{Bengtsson:1983pg}
  A.~K.~H.~Bengtsson, I.~Bengtsson and L.~Brink,
  %``Cubic Interaction Terms For Arbitrarily Extended Supermultiplets,''
  Nucl.\ Phys.\  B {\bf 227}, 41 (1983),
  %%CITATION = NUPHA,B227,41;%%
%\cite{Bengtsson:1983pd}
%\bibitem{Bengtsson:1983pd}
%  A.~K.~H.~Bengtsson, I.~Bengtsson and L.~Brink,
  %``Cubic Interaction Terms For Arbitrary Spin,''
  Nucl.\ Phys.\  B {\bf 227}, 31 (1983),
  %%CITATION = NUPHA,B227,31;%%
%\cite{Hussain:1988uk}
%\bibitem{Hussain:1988uk}
  F.~Hussain, G.~Thompson and P.~D.~Jarvis,
  %``Massive and Massless Gauge Fields of Any Spin and Symmetry,''
  Phys.\ Lett.\  B {\bf 216}, 139 (1989).
  %%CITATION = PHLTA,B216,139;%%


%\cite{Fronsdal:1978vb}
\bibitem{Fronsdal:1978vb}
  C.~Fronsdal,
  %``Singletons And Massless, Integral Spin Fields On De Sitter Space
  %(Elementary Particles In A Curved Space. 7,''
  Phys.\ Rev.\  D {\bf 20}, 848 (1979).
  %%CITATION = PHRVA,D20,848;%%




%\cite{Pashnev:1997rm}
\bibitem{Pashnev:1997rm}
  A.~Pashnev and M.~M.~Tsulaia,
  %``Dimensional reduction and BRST approach to the description of a Regge
  %trajectory,''
  Mod.\ Phys.\ Lett.\  A {\bf 12}, 861 (1997)
  [arXiv:hep-th/9703010].
  %%CITATION = MPLAE,A12,861;%%
%\cite{Buchbinder:2001bs}
%\bibitem{Buchbinder:2001bs}
   I.~L.~Buchbinder, A.~Pashnev and M.~Tsulaia,
   %``Lagrangian formulation of the massless higher integer spin fields in  the
   %AdS background,''
   Phys.\ Lett.\ B {\bf 523} (2001) 338
   [arXiv:hep-th/0109067].
   %%CITATION = HEP-TH 0109067;%%
   %\cite{Buchbinder:2002ry}
%%CITATION = HEP-TH 9803207;%%
%\bibitem{Buchbinder:2002ry}
 %  I.~L.~Buchbinder, A.~Pashnev and M.~Tsulaia,
   %``Massless higher spin fields in the AdS background and BRST  constructions
   %for nonlinear algebras,''
   [arXiv:hep-th/0206026].
   %%CITATION = HEP-TH 0206026;%%
%\cite{Buchbinder:2004gp}
%\bibitem{Buchbinder:2004gp}
   I.~L.~Buchbinder, V.~A.~Krykhtin and A.~Pashnev,
   %``BRST approach to Lagrangian construction for fermionic massless higher
   %spin fields,''
   Nucl.\ Phys.\ B {\bf 711}, 367 (2005)
   [arXiv:hep-th/0410215].
   %%CITATION = HEP-TH 0410215;%%
%\cite{Buchbinder:2006nu}
%\bibitem{Buchbinder:2006nu}
  I.~L.~Buchbinder, V.~A.~Krykhtin, L.~L.~Ryskina and H.~Takata,
  %``Gauge invariant Lagrangian construction for massive higher spin fermionic
  %fields,''
  Phys.\ Lett.\ B {\bf 641} (2006) 386
  [arXiv:hep-th/0603212].
  %%CITATION = HEP-TH 0603212;%%
%\cite{Buchbinder:2007ak}
%\bibitem{Buchbinder:2007ak}
  I.~L.~Buchbinder, A.~V.~Galajinsky and V.~A.~Krykhtin,
  %``Quartet unconstrained formulation for massless higher spin fields,''
  [arXiv:hep-th/0702161].
  %%CITATION = HEP-TH/0702161;%%
%\cite{Buchbinder:2007vq}
%\bibitem{Buchbinder:2007vq}
  I.~L.~Buchbinder, V.~A.~Krykhtin and A.~A.~Reshetnyak,
  %``BRST approach to Lagrangian construction for fermionic higher spin   fields
  %in AdS space,''
  [arXiv:hep-th/0703049].
  %%CITATION = HEP-TH/0703049;%%
%\cite{Buchbinder:2006ge}
%\bibitem{Buchbinder:2006ge}
  I.~L.~Buchbinder, V.~A.~Krykhtin and P.~M.~Lavrov,
  %``Gauge invariant Lagrangian formulation of higher spin massive bosonic field theory in AdS space,''
  Nucl.\ Phys.\  B {\bf 762} (2007) 344
  [arXiv:hep-th/0608005].
  %%CITATION = NUPHA,B762,344;%%
%\cite{Alfaro:2008ni}
%\bibitem{Alfaro:2008ni}
  J.~Alfaro and M.~Cambiaso,
  %``From isospin generators to BRST quantization of higher spin massless
  %fields,''
  [arXiv:0809.4298 [hep-th]].
  %%CITATION = ARXIV:0809.4298;%%
%\cite{Alkalaev:2009vm}
%\bibitem{Alkalaev:2009vm}
  K.~B.~Alkalaev and M.~Grigoriev,
  %``Unified BRST description of AdS gauge fields,''
  Nucl.\ Phys.\  B {\bf 835}, 197 (2010)
  [arXiv:0910.2690 ].
  %%CITATION = NUPHA,B835,197;%%
%\cite{Isaev:2008vi}
%\bibitem{Isaev:2008vi}
  A.~P.~Isaev, S.~O.~Krivonos and O.~V.~Ogievetsky,
  %``BRST operators for W algebras,''
  J.\ Math.\ Phys.\  {\bf 49}, 073512 (2008)
  [arXiv:0802.3781 [math-ph]].
  %%CITATION = JMAPA,49,073512;%%







%\cite{Metsaev:1997nj}
\bibitem{Metsaev:1997nj}
  R.~R.~Metsaev,
  %``Arbitrary spin massless bosonic fields in d-dimensional anti-de Sitter
  %space,''
 [ arXiv:hep-th/9810231],
  %%CITATION = HEP-TH/9810231;%%
%\cite{Metsaev:1998xg}
%\bibitem{Metsaev:1998xg}
%  R.~R.~Metsaev,
  %``Fermionic fields in the d-dimensional anti-de Sitter spacetime,''
  Phys.\ Lett.\  B {\bf 419}, 49 (1998)
  [arXiv:hep-th/9802097],
  %%CITATION = PHLTA,B419,49;%%
%\cite{Metsaev:2007rn}
%\bibitem{Metsaev:2007rn}
 % R.~R.~Metsaev,
  %``Cubic interaction vertices for fermionic and bosonic arbitrary spin
  %fields,''
  [arXiv:0712.3526 [hep-th]],
  %%CITATION = ARXIV:0712.3526;%%
%\cite{Metsaev:2005ar}
%\bibitem{Metsaev:2005ar}
 % R.~R.~Metsaev,
  %``Cubic interaction vertices for massive and massless higher spin fields,''
  Nucl.\ Phys.\  B {\bf 759}, 147 (2006)
  [arXiv:hep-th/0512342],
  %%CITATION = NUPHA,B759,147;%%


%\cite{Bekaert:2005jf}
\bibitem{Bekaert:2005jf}
   X.~Bekaert, N.~Boulanger and S.~Cnockaert,
   %``Spin three gauge theory revisited,''
   JHEP {\bf 0601}, 052 (2006)
   [arXiv:hep-th/0508048],
   %%CITATION = HEP-TH 0508048;%%
%\cite{Boulanger:2005br}
%\bibitem{Boulanger:2005br}
N.~Boulanger, S.~Leclercq and S.~Cnockaert,
%``Parity violating vertices for spin-3 gauge fields,''
Phys.\ Rev.\ D {\bf 73} (2006) 065019 [arXiv:hep-th/0509118],
%%CITATION = HEP-TH 0509118;%%
%\cite{Boulanger:2006gr}
%\bibitem{Boulanger:2006gr}
  N.~Boulanger and S.~Leclercq,
  %``Consistent couplings between spin-2 and spin-3 massless fields,''
  JHEP {\bf 0611} (2006) 034
  [arXiv:hep-th/0609221],
  %%CITATION = JHEPA,0611,034;%%
%\cite{Boulanger:2008tg}
%\bibitem{Boulanger:2008tg}
  N.~Boulanger, S.~Leclercq and P.~Sundell,
  %``On The Uniqueness of Minimal Coupling in Higher-Spin Gauge Theory,''
  JHEP {\bf 0808}, 056 (2008)
  [arXiv:0805.2764 [hep-th]].
  %%CITATION = JHEPA,0808,056;%%

%\cite{Zinoviev:2007js}
\bibitem{Zinoviev:2007js}
  Yu.~M.~Zinoviev,
  %``Massive N = 1 supermultiplets with arbitrary superspins,''
  Nucl.\ Phys.\  B {\bf 785}, 98 (2007)
  [arXiv:0704.1535 [hep-th]],
  %%CITATION = NUPHA,B785,98;%%
%\cite{Zinoviev:2008ck}
%\bibitem{Zinoviev:2008ck}
 % Yu.~M.~Zinoviev,
  %``On spin 3 interacting with gravity,''
  Class.\ Quant.\ Grav.\  {\bf 26}, 035022 (2009)
  [arXiv:0805.2226 [hep-th]],
  %%CITATION = CQGRD,26,035022;%%
%\cite{Zinoviev:2009wh}
%\bibitem{Zinoviev:2009wh}
 % Yu.~M.~Zinoviev,
  %``Note on antisymmetric spin-tensors,''
  JHEP {\bf 0904}, 035 (2009)
  [arXiv:0903.0262 [hep-th]],
  %%CITATION = JHEPA,0904,035;%%
%\cite{Zinoviev:2010cr}
%\bibitem{Zinoviev:2010cr}
 % Yu.~M.~Zinoviev,
  %``Spin 3 cubic vertices in a frame-like formalism,''
  [arXiv:1007.0158].
  %%CITATION = ARXIV:1007.0158;%%


%\cite{Manvelyan:2010jr}
\bibitem{Manvelyan:2010jr}
  R.~Manvelyan, K.~Mkrtchyan and W.~Ruehl,
  %``General trilinear interaction for arbitrary even higher spin gauge
  %fields,''
  Nucl.\ Phys.\  B {\bf 836}, 204 (2010)
  [arXiv:1003.2877],
  %%CITATION = NUPHA,B836,204;%%
%\cite{Manvelyan:2010wp}
%\bibitem{Manvelyan:2010wp}
%  R.~Manvelyan, K.~Mkrtchyan and W.~Ruehl,
  %``Direct construction of a cubic selfinteraction for higher spin gauge
  %fields,''
  [arXiv:1002.1358],
  %%CITATION = ARXIV:1002.1358;%%
%\cite{Manvelyan:2009vy}
%\bibitem{Manvelyan:2009vy}
 % R.~Manvelyan, K.~Mkrtchyan and W.~Ruhl,
  %``Off-shell construction of some trilinear higher spin gauge field
  %interactions,''
  Nucl.\ Phys.\  B {\bf 826}, 1 (2010)
  [arXiv:0903.0243 [hep-th]].
  %%CITATION = NUPHA,B826,1;%%
%\cite{Manvelyan:2009tf}
%\bibitem{Manvelyan:2009tf}
  R.~Manvelyan and K.~Mkrtchyan,
  %``Conformal invariant interaction of a scalar field with the higher spin
  %field in AdS_{D},''
  Mod.\ Phys.\ Lett.\  A {\bf 25}, 1333 (2010)
  [arXiv:0903.0058 [hep-th]].
  %%CITATION = MPLAE,A25,1333;%%




%\cite{Bastianelli:2009vj}
\bibitem{Bastianelli:2009vj}
  F.~Bastianelli and R.~Bonezzi,
  %``U(N) spinning particles and higher spin equations on complex manifolds,''
  JHEP {\bf 0903}, 063 (2009)
  [arXiv:0901.2311 [hep-th]].
  %%CITATION = JHEPA,0903,063;%%
%\cite{Bastianelli:2008nm}
%\bibitem{Bastianelli:2008nm}
  F.~Bastianelli, O.~Corradini and E.~Latini,
  %``Spinning particles and higher spin fields on (A)dS backgrounds,''
  JHEP {\bf 0811}, 054 (2008)
  [arXiv:0810.0188 [hep-th]].
  %%CITATION = JHEPA,0811,054;%%
%\cite{Corradini:2010ia}
%\bibitem{Corradini:2010ia}
  O.~Corradini,
  %``Half-integer Higher Spin Fields in (A)dS from Spinning Particle Models,''
  [arXiv:1006.4452],
  %%CITATION = ARXIV:1006.4452;%%
%\cite{Deser:2001us}
%\bibitem{Deser:2001us}
  S.~Deser and A.~Waldron,
  %``Partial masslessness of higher spins in (A)dS,''
  Nucl.\ Phys.\  B {\bf 607}, 577 (2001)
  [arXiv:hep-th/0103198].
  %%CITATION = NUPHA,B607,577;%%
%\cite{Gover:2008pt}
%\bibitem{Gover:2008pt}
  A.~R.~Gover, A.~Shaukat and A.~Waldron,
  %``Weyl Invariance and the Origins of Mass,''
  arXiv:0812.3364 [hep-th].
  %%CITATION = ARXIV:0812.3364;%%
%\cite{Cherney:2009md}
%\bibitem{Cherney:2009md}
  D.~Cherney, E.~Latini and A.~Waldron,
  %``Generalized Einstein Operator Generating Functions,''
  Phys.\ Lett.\  B {\bf 682}, 472 (2010)
  [arXiv:0909.4578 [Unknown]].
  %%CITATION = PHLTA,B682,472;%%
%\cite{Marnelius:2009uw}
%\bibitem{Marnelius:2009uw}
  R.~Marnelius,
  %``Lagrangian higher spin field theories from the O(N) extended supersymmetric
  %particle,''
  [arXiv:0906.2084 [hep-th]].
  %%CITATION = ARXIV:0906.2084;%%









%\cite{Sagnotti:2003qa}
\bibitem{Sagnotti:2003qa}
   A.~Sagnotti and M.~Tsulaia,
   %``On higher spins and the tensionless limit of string theory,''
   Nucl.\ Phys.\ B {\bf 682} (2004) 83
   [arXiv:hep-th/0311257].
   %%CITATION = HEP-TH 0311257;%%


%\cite{Buchbinder:2006eq}
\bibitem{Buchbinder:2006eq}
  I.~L.~Buchbinder, A.~Fotopoulos, A.~C.~Petkou and M.~Tsulaia,
  %``Constructing the cubic interaction vertex of higher spin gauge fields,''
  Phys.\ Rev.\  D {\bf 74} (2006) 105018
  [arXiv:hep-th/0609082].
  %%CITATION = PHRVA,D74,105018;%%

%\cite{Fotopoulos:2007nm}
\bibitem{Fotopoulos:2007nm}
  A.~Fotopoulos and M.~Tsulaia,
  %``Interacting Higher Spins and the High Energy Limit of the Bosonic String,''
  Phys.\ Rev.\  D {\bf 76}, 025014 (2007)
  [arXiv:0705.2939 [hep-th]].
  %%CITATION = PHRVA,D76,025014;%%


%\cite{Fotopoulos:2007yq}
\bibitem{Fotopoulos:2007yq}
  A.~Fotopoulos, N.~Irges, A.~C.~Petkou and M.~Tsulaia,
  %``Higher-Spin Gauge Fields Interacting with Scalars: The Lagrangian Cubic
  %Vertex,''
  JHEP {\bf 0710} (2007) 021
  [arXiv:0708.1399 [hep-th]].
  %%CITATION = JHEPA,0710,021;%%

%\cite{Fotopoulos:2009iw}
\bibitem{Fotopoulos:2009iw}
  A.~Fotopoulos and M.~Tsulaia,
  %``Current Exchanges for Reducible Higher Spin Multiplets and Gauge Fixing,''
  JHEP {\bf 0910}, 050 (2009)
  [arXiv:0907.4061 [hep-th]].
  %%CITATION = JHEPA,0910,050;%%


%\cite{Francia:2007qt}
\bibitem{Francia:2007qt}
  D.~Francia, J.~Mourad and A.~Sagnotti,
  %``Current exchanges and unconstrained higher spins,''
  Nucl.\ Phys.\  B {\bf 773}, 203 (2007)
  [arXiv:hep-th/0701163].
  %%CITATION = NUPHA,B773,203;%%
%\cite{Francia:2008hd}
%\bibitem{Francia:2008hd}
 % D.~Francia, J.~Mourad and A.~Sagnotti,
  %``(A)dS exchanges and partially-massless higher spins,''
  Nucl.\ Phys.\  B {\bf 804}, 383 (2008)
  [arXiv:0803.3832 [hep-th]].
  %%CITATION = NUPHA,B804,383;%%
%\bibitem{Sagnotti:2010jt}
  A.~Sagnotti,
  %``Higher Spins and Current Exchanges,''
  [arXiv:1002.3388].
  %%CITATION = ARXIV:1002.3388;%%

%\cite{Bekaert:2009ud}
\bibitem{Bekaert:2009ud}
  X.~Bekaert, E.~Joung and J.~Mourad,
  %``On higher spin interactions with matter,''
  JHEP {\bf 0905}, 126 (2009)
  [arXiv:0903.3338 [hep-th]].
  %%CITATION = JHEPA,0905,126;%%


%\cite{Francia:2002pt}
\bibitem{Francia:2002pt}
   D.~Francia and A.~Sagnotti,
   %``On the geometry of higher-spin gauge fields,''
   Class.\ Quant.\ Grav.\  {\bf 20} (2003) S473
   [arXiv:hep-th/0212185].
   %%CITATION = HEP-TH 0212185;%%
%\cite{Francia:2010qp}
%\bibitem{Francia:2010qp}
  D.~Francia,
  %``String theory triplets and higher-spin curvatures,''
  Phys.\ Lett.\  B {\bf 690}, 90 (2010)
  [arXiv:1001.5003].
  %%CITATION = PHLTA,B690,90;%%






%\cite{Sagnotti:2010at}
\bibitem{Sagnotti:2010at}
  A.~Sagnotti and M.~Taronna,
  %``String Lessons for Higher-Spin Interactions,''
  [arXiv:1006.5242],
  %%CITATION = ARXIV:1006.5242;%%
%\cite{Taronna:2010qq}
%\bibitem{Taronna:2010qq}
  M.~Taronna,
  %``Higher Spins and String Interactions,''
  [arXiv:1005.3061].
  %%CITATION = ARXIV:1005.3061;%%






%\cite{Polyakov:2010qs}
\bibitem{Polyakov:2010qs}
  D.~Polyakov,
  %``Gravitational Couplings of Higher Spins from String Theory,''
  [arXiv:1005.5512],
  %%CITATION = ARXIV:1005.5512;%%
%\cite{Polyakov:2009pk}
%\bibitem{Polyakov:2009pk}
 % D.~Polyakov,
  %``Interactions of Massless Higher Spin Fields From String Theory,''
  [arXiv:0910.5338].
  %%CITATION = ARXIV:0910.5338;%%


%\cite{Giombi:2009wh}
\bibitem{Giombi:2009wh}
  S.~Giombi and X.~Yin,
  %``Higher Spin Gauge Theory and Holography: The Three-Point Functions,''
  [arXiv:0912.3462],
  %%CITATION = ARXIV:0912.3462;%%
%\cite{Giombi:2010vg}
%\bibitem{Giombi:2010vg}
  %S.~Giombi and X.~Yin,
  %``Higher Spins in AdS and Twistorial Holography,''
  [arXiv:1004.3736].
  %%CITATION = ARXIV:1004.3736;%%







\end{thebibliography}

\documentstyle[12pt]{article}

\newcommand{\be}{\begin{equation}}
\newcommand{\bea}{\begin{eqnarray}}
\newcommand{\ee}{\end{equation}}
\newcommand{\eea}{\end{eqnarray}}

\def\theequation{\arabic{section}.\arabic{equation}}
\begin{document}
\setcounter{equation}{0}
 \renewcommand{\theequation}{\arabic{equation}}
\topmargin -1cm \oddsidemargin=0.25cm\evensidemargin=0.25cm
\setcounter{page}1
\renewcommand{\thefootnote}{\fnsymbol{footnote}}
\begin{flushright}
LTH 882
\end{flushright}
\vskip .7in
\begin{center}
{\Large \bf Current Exchanges for Reducible Higher Spin Modes on AdS
\footnote{Talk given at the XIXth International Colloquium on Integrable Systems and Quantum Symmetries,
Prague, Czech Republic, June 17-19, 2010} }
\vskip .7in {\large Angelos  Fotopoulos$^a$}\footnote{e-mail:
{\tt foto@to.infn.it}}
and {\large Mirian Tsulaia$^b$}\footnote{e-mail: {{\tt tsulaia@liv.ac.uk}, Associate member 
of the Centre of Particle Physics and Cosmology, Ilia State University, 0162, Tbilisi Georgia }} \vskip .2in
{$^a$\it Dipartimento di Fisica Teorica dell'Universit\`a di Torino
and INFN\\Sezione di Torino,
via P. Giuria 1, I-10125 Torino, Italy} \\
\vskip .2in {$^b$ \it Department of Mathematical Sciences, University of Liverpool,
Liverpool, L69 7ZL, United Kingdom}\\

\end{center}
\vskip .3in
\begin{abstract}
We show how to decompose a Lagrangian for reducible massless bosonic Higher Spin modes into the 
ones describing irreducible (Fronsdal) Higher Spin modes on a ${\cal D}$ dimensional $AdS$ space. Using this decomposition
we construct a new Nonabelian cubic interaction vertex  
for reducible higher spin modes and two scalars on $AdS$ from the already known vertex which involves irreducible (Fronsdal) modes.

\end{abstract}


Higher Spin gauge theories   (see \cite{Vasiliev:2004qz}--\cite{Fotopoulos:2008ka} for recent reviews) are usually formulated
either in frame--like \cite{Fradkin:1986qy}--\cite{Bandos:1999qf}  or  metric -- like \cite{Fronsdal:1978rb}--\cite{Francia:2002pt}
 approaches.

Recently several interesting cubic vertexes have been constructed in the metric -- like approach \cite{Metsaev:1997nj}, \cite{Bekaert:2005jf}
\cite{Zinoviev:2007js}, \cite{Manvelyan:2010jr}.
Bearing in mind a possible application of the reducible Higher Spin modes (described by the so called ``triplet'' \cite{Francia:2002pt} )
 for String Theory \cite{Sagnotti:2010at}--\cite{Polyakov:2010qs} and for AdS/CFT correspondence \cite{Giombi:2009wh}
we consider the problem of cubic interaction of a triplet on  AdS space. In particular we study the cubic
interaction of a triplet 
with two scalar fields.

The main result of this paper is twofold. Firstly, we show that the procedure derived in \cite{Fotopoulos:2009iw} to decompose
the free Lagrangian for reducible massless bosonic Higher Spin modes in a flat space time works
for an arbitrary dimensional AdS space as well. The second and more important result is that after 
this decomposition one can use the cubic vertex\footnote{Let us point out that the method given in \cite{Buchbinder:2006eq} 
describes construction of {\it nonabelian} cubic interaction vertexes, see also \cite{Fotopoulos:2007yq} for some explicit {\it nonabelian} examples.
A particular example of an abelian vertex given in \cite{Fotopoulos:2007nm} is in some sense a ``degenerate'' solution of the
method, where however the abelian property is maintained in a nontrivial way, due to the structure of the  ghost terms. }
of \cite{Fotopoulos:2007yq}, which describes an interaction of  irreducible (Fronsdal) Higher Spin modes
with two scalars, to obtain an interaction vertex for reducible Higher Spin modes with two scalars.
Obviously this technique can be applied not only for a particular vertex given in \cite{Fotopoulos:2007yq},
but for construction of more complicated interaction vertexes in AdS following a method given in \cite{Buchbinder:2006eq}.
The advantage of this approach is that a construction of interaction vertexes for triplets
in AdS is often technically complicated due to repeated commutators between covariant derivatives 
and the double tracelessness condition for irreducible Higher Spin modes makes the problem at hand 
considerably simpler.

Let us start from a free Lagrangian describing the propagation of reducible massless Higher Spin modes
on a ${\cal D}$ dimensional AdS space.
It contains a field $\varphi_{\mu_1,...,\mu_s}(x)$ of the rank $s$, a field $C_{\mu_1,...,\mu_{s-1}}(x)$ of  rank $s-1$ and 
a field $D_{\mu_1,...,\mu_{s-2}}(x)$ of a rank $s-2$ and has the form \cite{Sagnotti:2003qa} (see also \cite{Fotopoulos:2008ka}
for the details of the construction),
\begin{eqnarray} \nonumber
{\cal L} &=& - \, \frac{1}{2}\ (\nabla_\mu \varphi)^2 \ + \ s\,
\nabla \cdot \varphi \, C \ + \ s(s-1)\, \nabla \cdot C \, D \
 + \ \frac{s(s-1)}{2} \, (\nabla_\mu D)^2 \ - \ \frac{s}{2} \,
C^2 \\ \nonumber &+& \ \frac{s(s-1)}{2L^2}\, {(\varphi^{'})}^2 \ - \
\frac{s(s-1)(s-2)(s-3)}{2L^2} \, {(D^{'})}^2
  \ - \ \frac{4s(s-1)}{L^2} \, D \, \varphi^{'} \\
&-& \ \frac{1}{2L^2} \, \left[ (s-2)({\cal D}+s-3) \, - \, s
\right] {\varphi}^2 \ + \ \frac{s(s-1)}{2L^2} \, \left[ s({\cal
D}+s-2)+6 \right]\,  D^2  , \label{LtripletBADS}
\end{eqnarray}
The
symbol $\nabla \cdot$ means divergence, while $\nabla$  is
symmetrized action of $\nabla_\mu$ on a tensor. The symbol
${}^\prime$ means that we take the trace of a field.
Multiplication of a tensor by the metric $g$ implies 
symmetrized multiplication, i.e., if $A$ is a vector $A_\mu$ we
have $g A = g_{( \mu \nu} A_{\rho)}= g_{\mu \nu} A_\rho +g_{\mu \rho} A_\nu+g_{\nu \rho}
A_\mu $.
This Lagrangian is invariant under the gauge transformations
with  parameter  $\Lambda_{\mu_1,...,\mu_{s-1}}(x)$
\begin{eqnarray}
&& \delta \varphi \ = \ \nabla \, \Lambda \ , \nonumber \\
&& \delta  C  \ = \ \Box \;
\Lambda + \frac{(s-1)(3-s-{\cal
D})}{L^2}\
  \Lambda
+ \frac{2}{L^2} \,  g \; \Lambda^\prime
 \nonumber \\
&& \delta D \ = \ \nabla \cdot
\Lambda \ . \label{adstripletgauge}
\end{eqnarray}
Let us note that the field $C(x)$ has no kinetic term and  can be eliminated via its own equations
of motion to obtain
\begin{eqnarray}\label{Lcomp2}
{\cal L} &=& - \, \frac{1}{2}\ (\nabla_\mu \varphi)^2 \ + \
\frac{s}{2} \,
(\nabla \cdot \varphi)^2 \ + \ s(s-1)\, \nabla \cdot \nabla \cdot \varphi \, D \ \nonumber \\
&+& \ s(s-1) \ (\nabla_\mu D)^2 \ + \ \frac{s(s-1)(s-2)}{2} \,
(\nabla \cdot D)^2  \\ \nonumber
&+& \frac{s(s-1)}{2L^2}\, {(\varphi^{'})}^2 \ - \
\frac{s(s-1)(s-2)(s-3)}{2L^2} \, {(D^{'})}^2
  \ - \ \frac{4s(s-1)}{L^2} \, D \, \varphi^{'} \\ \nonumber
&-& \ \frac{1}{2L^2} \, \left[ (s-2)({\cal D}+s-3) \, - \, s
\right] {\varphi}^2 \ + \ \frac{s(s-1)}{2L^2} \, \left[ s({\cal
D}+s-2)+6 \right]\,  D^2  .
\end{eqnarray}
Now we would like to decompose this Lagrangian in terms of irreducible (Fronsdal) \cite{Fronsdal:1978vb}
modes, following the procedure given in \cite{Fotopoulos:2009iw} for a  Minkowski space.

Let us start with the simplest example of a $s=2$ triplet which contains fields $\varphi_{\mu \nu}(x)$,
$C_\mu(x)$ and $D(x)$. Let us make the ansatz 
\begin{equation}
\varphi_{\mu \nu}= \Psi_{\mu \nu} + \frac{1}{{\cal D} -2} g_{\mu \nu} \Psi, \quad 
 \varphi^{'}-2D=\Psi.
\end{equation}
Inserting these expressions back to the Lagrangian (\ref{Lcomp2}) for $s=2$
one obtains
 \begin{eqnarray}\label{Ls2} \nonumber
 {\cal L}&=& -{1 \over 2} (\nabla_\mu \Psi_{\rho \sigma})^2 +
 (\nabla_\nu \Psi^\nu_\mu)^2 + \Psi^{'}
 \nabla_{\mu}\partial_{\nu} \Psi^{\mu \nu} + {1\over 2}
 (\nabla_\mu \Psi^{'})^2 - {1 \over 2({\cal D}-2)} (\nabla_\mu \Psi)^2 \\ 
&+& \frac{1}{L^2} (\Psi_{\mu \nu})^2 + \frac{{\cal D}-3}{L^2 ({\cal D}-2)} (\Psi)^2 + \frac{{\cal D}-3}{2L^2 } {(\Psi^{'})}^2
 \end{eqnarray}
Therefore, the initial Lagrangian (\ref{Lcomp2}) has been decomposed into a sum of two Fronsdal Lagrangians
for $s=2$ field $\Psi_{\mu \nu}$ with the gauge transformation law $\delta \Psi_{\mu \nu} = \nabla_\mu \Lambda_\nu +\nabla_\mu \Lambda_\nu $
 and a gauge invariant scalar $\Psi$.

Let us describe this procedure for the spin $4$ triplet, since in this case both a constraint on the parameter of gauge transformations
and an off--shell constraint on the gauge field arise. Let us use the substitution \cite{Fotopoulos:2009iw}
\begin{eqnarray}\label{phiDs4}
\varphi^{(4)} &=& \Psi^{(4)} + {1\over {\cal D}+2} g \Psi^{(2)} + {1\over {\cal D}({\cal D}-2)} {(g)}^2 \Psi^{(0)} \nonumber \\
D&=& {1\over 2} [ {\Psi'}^{(4)} +{2\over {\cal D}+2} \Psi^{(2)} + {1\over {\cal D}+2} g
{\Psi'}^{(2)} +{2\over {\cal D}({\cal D}-2)} g \Psi^{(0)}].
\end{eqnarray}
The field $\Psi^{(4)}$ is doubly traceless and transforms under the gauge transformations as
\begin{equation}
\delta \Psi^{(4)}= \nabla \tilde \Lambda,
 \quad \tilde \Lambda= \Lambda - {1\over {\cal D}+2}
 \eta \Lambda' 
\end{equation}
Inserting these expressions into the Lagrangian  (\ref{Lcomp2}) one can see again that it decomposes into the sum 
of Fronsdal modes with spins $4,2$ and $0$, described by the fields $ \Psi^{(4)},  \Psi^{(2)}$ and  $\Psi^{(0)}$.
 
One can further generalize this procedure for an arbitrary spin. In particular take
\begin{eqnarray}\label{fD}
\varphi &=&  \sum_{k=0}^{[{s \over 2}]}
\tilde{\rho}_k({\cal D},s) {(g)}^k \Psi^{(s-2k)} \nonumber \\
D &=& {1\over
2}\sum_{k=0}^{[{s \over 2}]-1} \tilde{\rho}_k({\cal D},s) {(g)}^k
\Psi^{'({s-2k})} + \sum_{k=1}^{[{s \over 2}]} \tilde{\rho}_k({\cal D},s)
{(g)}^{k-1} \Psi^{(s-2k)}.
\end{eqnarray}
and 
\begin{equation}\label{tLambda}
{\tilde \Lambda}_{s-1-2k}= \sum_{q=0}^{[{s\over 2}]}
\rho_q({\cal D},s-2k-1) {(g)}^q \Lambda^{[q+k]({s-1})}
\end{equation}
with
\begin{equation}\label{rho}
\rho_q({\cal D}-2,s)=  {(-1)^q
({\cal D}+2(s-q-3))!! \over ({\cal D}+2(s-3))!!}, \quad \tilde{\rho}_k({\cal D},s)= {({\cal D}+2(s-2k-2))!! \over ({\cal D}+2(s-k-2))!!}
\end{equation}
and $[q+k]$ denotes a number of traces.
Finally, one can show that the normalization factor
for propagators for each of individual Fronsdal mode
i.e., the
inverse of the prefactor of  $(\nabla_\mu \Psi^{(s-2k)})^2$ terms
multiplied by 2 is
\begin{equation}\label{Q}
Q(s,k,{\cal D})= { 2^k k! (s-2k)!\over s!{\tilde \rho}_k({\cal D},s)}.
\end{equation}
Now let us build a cubic interaction vertex of a Higher Spin triplet with two scalars on AdS. To this end let us use
the corresponding vertex for an  individual Fronsdal mode \cite{Fotopoulos:2007yq}
\begin{equation} \label{INV}
{\cal L}_{int}^{00s}= \Psi^{(s)} \cdot J_s + [\frac{s-1}{6L^2} [2s^2 + (3 {\cal D}-4)s -6] - \frac{s-2}{L^2}]\Psi^{' (s)} \cdot J_{s-2} 
\end{equation}
where
\begin{equation}
J^{1;2}_{s-2q}=\sum_{r=0}^{s-2q}  C_{s-2q}^{r}  (-1)^r 
(\nabla^{\mu_1} \dots
\nabla^{\mu_r}
\phi_1)\ (\nabla^{\mu_{r+1}} \dots
\nabla^{\mu_{s-2q}}\phi_2)\
\end{equation}
Therefore  multiplying    interacting vertexes (\ref{INV}) with the appropriate factor
(\ref{Q}) and adding them to the free Lagrangian (\ref{Lcomp2}), one finds the expression for a  cubic Lagrangian
describing the interaction of reducible Higher Spin modes with two scalars on AdS.
After that one can perform a current- current exchange procedure following the lines of
\cite{Francia:2007qt}, \cite{Bekaert:2009ud}, \cite{Fotopoulos:2009iw}.

\noindent {\bf Acknowledgements.}
It is a pleasure to thank A.P. Isaev, S.O. Krivonos and A.O. Sutulin
for valuable discussions.
The work of A. F. was supported by
an INFN postdoctoral fellowship and partly supported by the Italian
MIUR-PRIN
contract 20075ATT78.
The work of M.T. has been supported by a STFC rolling grant ST/G00062X/1.

\end{document}